OpenStudy (anonymous):

Find the area of the surface obtained by rotating y = cosh x over [−ln 2, ln 2] around the x–axis.

OpenStudy (accessdenied):

Do you know the integration surface area formula for this case?

OpenStudy (accessdenied):

It is very similar to taking the circumference of a circle with the function as your radius, and multiplying by a tiny width along the curve called ds / adding up all those little areas.

OpenStudy (accessdenied):

We generally change out ds for dx or dy using the definition given to ds:  Factoring (dx)^2 and taking the square root of both sides: (ds)^2 = (dx)^2 + (dy)^2 (ds)^2 = (1 + (dy/dx)^2 ) (dx)^2 ds = sqrt(1 + dy/dx ^2) dx

OpenStudy (accessdenied):

This is a good resource for more information if I don't return for any responses. :) http://tutorial.math.lamar.edu/Classes/CalcII/SurfaceArea.aspx